A circular tank with a diameter at the bottom of D 20 m and with vertical walls been filled with a flow of 240 m3/h. what will be the increase of the tank depth after 15 minutes, assuming a constant flow during this period of time?
1
Expert's answer
2020-04-23T13:13:47-0400

Volume of tank="\pi r^2h ," h= depth of tank and the dia of tank is =20m so radius of tank is 10 m .

Now, flow rate is constant =Q= 240 "\frac{m^3}{h}"

and flow is for 15 min means "times= \frac{1}{4} \times 240 = 60 m^3"

SO, change in volume= dV

"\pi r^2 (h_2-h_1)= 60"

change in depth or increase in depth= "\frac{60}{3.14\times 10\times 10}" = 0.19 m